# HistDiST: Histopathological Diffusion-based Stain Transfer

### v1

@@ 3.0.5 Ablation Studies:
$\eta$ constant 0.4500 0.598 41.7
$\eta$ cosine 0.4426 0.559 34.3
$\eta$ cosine start 20% 0.4360 0.552 31.3
$\eta$ cosine start 30% 0.4328 0.547 30.1
@@
Figure 3: ER translation outputs across methods (see Table 2 for abbreviations).
@@
To address this, we conduct joint training on H&E generation and H&E-to-IHC translation, and at inference apply DDIM inversion followed by DDIM sampling with $\eta$ cosine start 20% noise schedule, which greatly enhances structural preservation. Finally, incorporating Phikon-extracted morphological embeddings into U-Net’s cross-attention layers further refines feature-space alignment and enhances structural integrity and molecular fidelilty, yielding the best overall performance. This demonstrates that explicit pathology-aware conditioning is critical for accurate molecular reconstruction and high-fidelity stain translation. In most of the cases, MDCL often introduces artifacts or fails to capture fine molecular details, whereas HistDiST better preserves molecular expression patterns and produces more precise IHC intensity distributions, aligning closely with GT IHC.
@@
4 Conclusion
We present HistDiST, an LDM-based framework for H&E-to-IHC stain translation, integrating pathology-aware conditioning, DDIM inversion, and zero terminal SNR enforcement to enhance structural fidelity and molecular accuracy. We also introduce Molecular Retrieval Accuracy (MRA), a novel metric for pathology-aware evaluation. Experiments on MIST and BCI datasets demonstrate that HistDiST achieves superior visual fidelity, molecular accuracy and structural preservation compared to the existing approaches.

### v2

@@ 4.0.5 Ablation Studies and Qualitative Results:
$\eta$ constant 0.450 0.598 41.7
$\eta$ cosine 0.443 0.559 34.3
$\eta$ cosine start 20% 0.436 0.552 31.3
$\eta$ cosine start 30% 0.433 0.547 30.1
@@
Figure 3: ER translation outputs across methods (see Table 2 for abbreviations). The first two columns show the H&E input and its corresponding ground-truth ER pair from the dataset, the next two columns (MDCL and ControlNet) show the previous top-performing GAN- and Diffusion-based approaches, and the last four columns are the incremental ablation of the components of our model, with our final model shown as the last column.
@@
To address this, we conduct joint training on H&E generation and H&E-to-IHC translation, and at inference, apply DDIM inversion followed by DDIM sampling with $\eta$ cosine start 20% noise schedule, which greatly enhances structural preservation. Finally, incorporating Phikon-extracted morphological embeddings into U-Net’s cross-attention layers further refines feature-space alignment and enhances structural integrity and molecular fidelity, yielding the best overall performance. This demonstrates that explicit pathology-aware conditioning is critical for accurate molecular reconstruction and high-fidelity stain translation. The qualitative comparison of our method against the best-performing GAN- and Diffusion-based approaches, MDCL and ControlNet, is also shown in Fig. 3. While ControlNet mostly provides improvements over the InstructPix2Pix baseline, it fails to preserve the structure accurately since acquiring Canny edges for low-contrast H&E images is challenging, and also fails to obtain accurate IHC expressions. Furthermore, in most cases, MDCL often introduces artifacts or fails to capture fine molecular details. On the other hand, HistDiST better preserves molecular expression patterns and produces more precise IHC intensity distributions, aligning closely with GT IHC. For instance, in the fourth row of Fig. 3, MDCL and ControlNet incorrectly predict the molecular expression pattern (shown with a high expression level) while our method accurately predicts the patterns as low expression levels (last column).
@@
5 Conclusion and Limitations
We present HistDiST, an LDM-based framework for H&E-to-IHC stain translation, integrating pathology-aware conditioning, DDIM inversion, and zero terminal SNR enforcement to enhance structural fidelity and molecular accuracy. Notably, the inversion process helps to preserve the structure; still, we employ controlled stochasticity during denoising as fully-deterministic sampling negatively affects the fidelity of generated images. We also introduce Molecular Retrieval Accuracy (MRA), a novel metric for pathology-aware evaluation based on the alignment of pathological features. Experiments on MIST and BCI datasets demonstrate that HistDiST achieves superior visual fidelity, molecular accuracy and structural preservation compared to the existing approaches. Nevertheless, we emphasize that this method is not clinically applicable at this stage; additional validation and analysis (e.g., expert review or downstream task performance) would be required before considering diagnostic use. Future studies can further explore the potential of stain transfer methods by benchmarking their utility in downstream tasks such as patch-level classification or segmentation.